OpenStudy (anonymous):

Diff eq question! (round 3...)

OpenStudy (anonymous):

\[\Gamma(p + 1) = \int\limits_{0}^{\infty} e ^{-x} x ^{p}dx\] Show that, for p > 0,\[\Gamma(p+1) = p \Gamma(p)\]

OpenStudy (anonymous):

You can start by writing what \[ p\Gamma(p) \] would look like, based on the definition above.

OpenStudy (anonymous):

Wellp, I'm bored, here ya go: \[ p\Gamma(p) = p \cdot \int_0^\infty e^{-x}x^{p-1} dx \] \[ = \int_0^\infty e^{-x} \left( p x^{p-1} dx \right) \] Now we integrate by parts to get: \[ =\left. e^{-x}\cdot x^p \right|_0^\infty - \int_0^\infty\left(- e^{-x}\right) x^p dx \] \[ = 0 + \int_0^\infty e^{-x} x^p dx = \Gamma(p+1) \]